A proton of mass 1.67*10^-27 kg is accelerated by a field of 600 N/C. How fast in m/s is it travelling after 20 nanoseconds if it has a charge of 1.60*10^-19 C?
Answer #1

electric force on proton, F = q*E

m*a = q*E

a = q*E/m

= 1.6*10^-19*600/(1.67*10^-27)

= 5.75*10^10 m/s^2

Apply, v = u + a*t

= 0 + 5.75*10^10*20*10^-9

= 1.15*10^3 m/s